Solve each problem involving consecutive integers.
Find three consecutive integers such that the sum of the first and twice the second is 17 more than twice the third.
I know that consecutive integers are x, x+1, x+2 and they are added together, but I don't know how to set up the problem.
So the first number would be x
The second number would be x+1
And the third number would be x+2
Or at least that's what I understand and have been thinking.

Answer by ankor@dixie-net.com(22740)

Solve each problem involving consecutive integers.
Find three consecutive integers such that the sum of the first and twice the second is 17 more than twice the third.
I know that consecutive integers are x, x+1, x+2 and they are added together, but I don't know how to set up the problem.
So the first number would be x
The second number would be x+1
And the third number would be x+2
:
that's right then just write an equation for the statement:
"three consecutive integers such that the sum of the first and twice the second is 17 more than twice the third."
x + 2(x+1) = 2(x+2) + 17
x + 2x + 2 = 2x + 4 + 17
3x + 2 = 2x + 21
3x - 2x = 21 - 2
You get the idea now I am sure